Astronomy Report – January 4th

Crew 202 Astronomy Report 04-Jan-2019
Crew Astronomer: Cesare Guariniello

MDRS ROBOTIC OBSERVATORY
Robotic Telescope: MDRS-WF
Objects Viewed: M66 and Horsehead Nebula (photos from yesterday’s observation)

Problems Encountered: none

MUSK OBSERVATORY
Solar Features Observed: There were no interesting solar features. The astronomer worked on focus and tuning with both the eyepiece and the camera

Problems Encountered: none

Journalist Report – January 4th

MDRS Crew 202, Journalist Report

Sol 6 – 01/04/2018

Name the space movie (or show) given the following quote. Answer at the end of the Report:

I sometimes catch myself looking up at the Moon, remembering the changes of fortune in our long voyage, thinking of the thousands of people who worked to bring the three of us home. I look up at the Moon and wonder, when will we be going back, and who will that be?

Our morning went from 0 to 100 when we chose to start the day with an ab workout. The Engineer and Executive Officer have been betting on whether the other could hold a 5 minute plank which has evolved to various competitions of exercises. While their physical prowess is impressive, the aches and pains of my sore bones are looking forward to leading a meditation tomorrow instead of full blown P90X.

The Geologist (GEO), Health & Safety Officer (HSO), and Commander (CMD) went on the 5th Extra-Vehicular Activity (EVA) of our mission while the Executive Officer (EXO), Engineer (ENG), and I (Crew Journalist = JOU) stayed behind to work through our own projects and relax after going on the previous two EVAs. This was a big day for GEO and HSO since they haven’t left the Habitat in three days. Three days of living in a two-story Corn Silo. The only relief coming from the sun shining through the four windows in the living room, all no larger than a beach ball. Needless to say, they were excited to pretend to breathe the fresh air through their suits and touch the ground through their heavily padded hands and feet.

It is not the large parts of living on Earth you miss, such as the scenery or restaurants. It is the little things you miss as your brain teases you with cravings and wishes. An innocent thought popping up like “ice cream sounds nice”, and then quickly realizing there’s no possible way to concoct anything even related to ice cream from the dehydrated, shelf stable dust we’ve been given to live off of…. Or standing in a hot shower. Steam filling the room. The water gently rolling down your face without using a wash cloth to place it there…

Ok. I’m back. Daydreaming for a minute about that shower. And ice cream. Maybe both at the same time. It’s difficult to plan what to indulge in first once we get back to Earth. As a gentle reminder to the reader, myself, and the crew, it’s our 7th day on Mars, nearly a full week after leaving Earth. We ran out of fruit today. The canned spam is taunting us from the cupboards. Let us hope we never reach that point of desperation.

Speaking of food, our CMD and GEO have never had ramen. You know, the $0.29 plastic packaged sodium bomb that you lived off of in college because you blew what little money you had on Insomniac cookies the night before. So… after a long day of trekking in the Martian snow, boots soiled, and noses red, our EVA crew returned to nice warm pot of noodles in bouillon cubes. Their verdict was “It was better than I expected”. That’s essentially a 10/10 review for ramen.

The second floor of the habitat is always filled with laughter. It’s a good thing we live on a desolate planet because we would wake the whole neighborhood with the joy, conversation, and howling echoing through the walls. Even through the dehydrated food dirt and dreaming of creamier bites and cleaner days, the overall mood of Crew 202 is bright. It’s another wonderful day on the Red Planet as we enter day 2 without a major crisis. Yes, I did knock on the wood planked floors before sending this journal.

Movie (or Show) Answer: Apollo 13

Greenhab Report – January 4th

Hi Mission Support,

Crew 202 Greenhab Report 04-Jan-2019
Greenhab Officer:
Jake Qiu

Environmental Control:
Heating
Cooling w/ ambient air (1hrs)

40% Shade Cloth on
80% Shade Cloth on

Average Temperatures:
Low: 23.2°C
High: 39.5°C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
15 gallons

Water in Blue Tank (lbs):
81%

Times of Watering for Crops:
1014
1855

Changes to crops:
* Some of the microgreens showing good progress
* Lettuce Red leafs showing good progress
* Some carrots still lacking progress
* Cress displaying good progress
* One microgreen tray for crew harvest was wrapped in seran wrap to trap humidity
* Cucumbers starting to look worse

Narrative:
Watered plants throughout the day as needed. Boiled water and allowed to cool for
microgreens.

Experiment Results:
Date: SOL6 Crew 202 (SOL14 overall)

Watered all 21 trays as needed

Watered previous crew 201 experiments

Harvest:

Support/Supplies Needed:

Best Regards,

Jake

EVA Report – January 4th

Hi Mission Support,

Crew 202 EVA Report 04-JAN-2019

EVA #5

Author: Ellen Czaplinski

Purpose of EVA: To collect various geological samples and measure
radiation in several locations

Start time: 10:42

End time: 1:40

Narrative: Today’s EVA had a unique start, as our crew journalist
conducted a “20 questions” interview with the crew geologist while we
were putting on flight suits and heading out to the rovers. Our first
destination was Greenstone Road, which was a great place to collect
sand samples for our geology research project. This area was full of
loose sand, ripples, and sandstone. Two samples were collected here,
and we also took their spectra in situ using the TREK portable
spectrometer. We also took radiation measurements of this environment.

The next stop was SE of Kissing Camel Range (between Greenstone Rd and
Kissing Camel Range), north of our first stop. No geological samples
were taken here, but our crew commander took radiation measurements.

Our last stop of the EVA was on the north side of Kissing Camel Range,
west of Cow Dung Road. Using the rock hammer, we knocked off parts of
the Quaternary-aged boulders that had fallen from the top of the
range. Parts on the inside of these rocks resembled green, volcanic
minerals, similar to ones collected from EVA #4. The field spectra of
this sample was indeed different from the clays and shales nearby.
Further west at this site, we walked up a dried gulley and collected
three more samples here. Measuring these samples’ spectra revealed
more shale and gypsum combined with another mineral that will be
determined through more analysis of the spectra.

Destination: Greenstone Road, Kissing Camel Range

Coordinates (use UTM NAD27 CONUS): E520200, N4248700

Participants: Cesare Guariniello (CMD), Ellen Czaplinski (GEO), Jake Qiu (HSO)

Road(s) and routes per MDRS Map: Drive on Cow Dung Road, then walk
along Greenstone road

Mode of travel: Curiosity and Spirit rovers, walking

Best regards and hope you’re having a good time being assistant director!

Operations Report – January 4th

Howdy Mission Support!

I am about to run out to the ScienceDome to check the SOC

Crew 202 Operations Report 04-JAN-2019

SOL: 5

Name of person filing report: Kasey Hilton

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run): 16hr 29min; Turned on last night (03Jan2019) at 16:36; Turned off this morning (04Jan2019) at 9:05; Turned on tonight (04Jan2019) at 15:02

Solar SOC – Turned on (03Jan2019) 83%; Turned off (04Jan2019) 100%; Turned on (04Jan2019) 81%

Diesel Reading – 73%

Propane Reading – 48%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 75%; 400 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 75%; 230 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:06 to refill tank

Water Meter: 01398024 units

Toilet tank emptied: Yes, still very stinky ://

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 64.7

Beginning charge: 100%

Ending charge: 44%

Currently charging: Yes

Opportunity rover used: No; Still out of brake fluid

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 62.9

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Today was the first day that nothing had broken!

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Best,

Kasey

Sol Summary – January 4th

Crew 202 Sol Summary Report 04-Jan-2019

Sol: 6

Summary Title: Settling down

Author’s name: Cesare Guariniello

Mission Status: All projects are moving forward, with some ups and downs. We are collecting plenty of data, and each of us is giving much contribution, both with our own research projects and as test subjects for our crewmates.

Sol Activity Summary: Our morning routine continues to be a varying mix of useful stretching and conditioning exercises. Today the crew members who stayed in the habitat yesterday went on EVA and vice versa. The physical and mental stress of the mission is beginning to creep through, but the crew is reacting very well. Time is flying, and it is already Sol 6 of our mission, we now feel this is "our" place, that we live in and that we take care of. It is a great view when coming back from EVA. The crew keeps having a great time together: before each EVA, the whole crew is involved in the preparation, and those who stay in the habitat are always ready to welcome the EVA crew with some hot meal. After the usual afternoon of rest and work on the projects, the crew is preparing for the evening (yesterday we had a movie night).

Look Ahead Plan: Tomorrow we will have a geology EVA towards canyons, pending advice on practicability of roads.

Anomalies in work: None

Weather: Still cold but getting warmer. Some snow on the ground and on the north face of hills.

Crew Physical Status: Exercises and much work are giving us health and some soreness. Morale is high!

EVA: The Commander, Health and Safety Officer, and Crew Geologist had an EVA to Greenstone Rd and to the region North of Kissing Camel Range, to collect sand and shale samples.

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, EVA report, EVA request, Astronomy Report, Journalist Report

Support Requested: None

Cesare
Commander, MDRS 202